# Defining Neural Network Architecture through Polytope Structures of Dataset

### v1

@@ 5 Conclusion
In this paper, we investigated the geometric characteristics of datasets and neural network architecture. Specifically, we established both lower and upper bounds on the necessary widths of network architectures for classifying given data manifolds, relying on its polytope-basis cover. Furthermore, we extended these insights to simplicial complexes or spaces consisting of prismatic polytopes, shedding light on how the width bound varies in response to the geometric complexity of the dataset. Conversely, we also demonstrated that the polytope structure of datasets can be inspected by training neural networks. We proposed such an algorithm, and our experimental results unveiled that each class within MNIST, Fashion-MNIST, and CIFAR10 datasets can be distiguished by at most two polytopes, with each comprising less than 30 faces. We further analyzed that the number of faces in the polytope serves as an indicator of the geometric complexity of each class. Our empirical investigations unveil that neural network indeed converges to a polytope-basis cover of dataset, and conversely, it is possible to inspect geometrical features of the dataset by training neural networks. Our findings for MNIST, Fashion-MNIST, and CIFAR10 datasets not only underscore the inherent simplicity of the geometry within these real-world datasets but also shed light on the investigating geometric complexity of individual classes.
@@
The optimality of the polytope-basis cover obtained by Algorithm 1 has not been clarified, which remains an interesting avenue for future research.
@@ Appendix A. Related works
Appendix B. Geometric analysis on the manifold of real-world datasets
@@ C.2 A polytope-basis cover derived from a trained two-layer ReLU network
              C.3 The efficient algorithm to find real-world datasets
              C.4 Comparison of proposed algorithms.
@@ Appendix A Related Works
On the other hand, there has been a burgeoning interest in investigating polytope structures induced by deep ReLU networks (Fawzi et al., 2018; Xu et al., 2021; Alfarra et al., 2022; Vincent & Schwager, 2022; Black et al., 2022; Haase et al., 2023; Liu et al., 2023; Fan et al., 2023; Huchette et al., 2023; Vallin et al., 2023; Piwek et al., 2023). Masden (2022) introduced algorithms capable of extracting the polytope structure inherent in networks and deriving topological properties of the decision boundary. Carlsson (2019) and Vallin et al. (2023) considered the pre-image of ReLU networks, characterizing the geometric shapes of the decision boundary to gain an understanding of the polytope partitions of deep ReLU networks.
@@
Recent years have witnessed a surge in research dedicated to unraveling the convexity of ReLU networks concerning their input. Crucially, it has been established that the ReLU activation function demonstrates convexity concerning its input when composited weights are all positive (Proposition 1 in (Amos et al., 2017), cf. Lemma F.4). This discovery led to the inception of Input Convex Neural Networks (ICNNs) by Amos et al. (2017), a characteristic that has been effectively leveraged in various applications (Makkuva et al., 2020; Chen et al., 2020; Fan et al., 2020; Bünning et al., 2021; Balestriero et al., 2022). Notably, Balestriero et al. (2022) applied this idea and proposed DeepHull, the algorithm to approximate the convex hull by convex deep networks. On the other hand, Sivaprasad et al. (2021) asserted that these convex neural networks exhibit self-regularization effects, demonstrating superior generalization performance in specific tasks. Additionally, research has delved into representing neural networks as the difference of convex (DC) functions (Sankaranarayanan & Rengaswamy, 2022; Piwek et al., 2023; Beltran-Royo et al., 2024). Sankaranarayanan & Rengaswamy (2022) employed Linear Programming to optimize polyhedral DC functions, capitalizing on the piecewise linearity induced by the ReLU activation function.
@@
Several studies have delved into the intricate relationship between the geometric features of datasets and neural network training, often referred to as the multiple manifold problem (Goldt et al., 2020; Buchanan et al., 2020; Wang et al., 2021; Chen et al., 2022; Tiwari & Konidaris, 2022). This problem typically involves the binary classification of two low-dimensional manifolds by neural networks. For instance, Buchanan et al. (2020) and Wang et al. (2021) focused on the task of distinguishing between two curves (i.e., one-dimensional manifolds), investigating the convergence speed and generalization concerning the geometric features of the dataset. Similarly, in the context of low-dimensional data manifolds, Tiwari & Konidaris (2022) examined the effects of data geometry on the complexity of trained neural networks, measuring the distance to the manifold. In a related vein, Dirksen et al. (2022) considered the separation problem with randomly initialized ReLU networks, explicitly linking required widths to weight initialization. Additionally, numerous empirical studies have supported the implicit relationship between network architecture and the geometric complexity or topological structure of the data manifold (Fawzi et al., 2018; Kim et al., 2020; Cohen et al., 2020; Birdal et al., 2021; Barannikov et al., 2021a, b; Naitzat et al., 2020; Hajij & Istvan, 2020, 2021; Magai & Ayzenberg, 2022). These theoretical and empirical findings suggest a high correlation between neural network architecture and the training dataset: a more complicated dataset requires a more complex architecture.
However, a notable gap still exists in the literature when it comes to explicitly constructing a neural network in practice. For example, as introduced in Section 1, it remains unknown which architecture of neural networks can or cannot completely classify a given dataset with explicit construcution of neural networks.
@@ Appendix B Geometric simplicity of real-world datasets
In this section, we present empirical results illustrating how the number of faces of covering polytopes reflects the geometric characteristics of a class. Specifically, we examine a classification task under varying levels of noisy labels. We focus on the class $\{1\}$ in each dataset and manipulate the noise levels ($r$) across values of 0, 0.01, 0.1, 0.25, 0.5, and 0.90 while maintaining the total number of data points in the class. The noisy class with noise level $r$ is denoted by $\{1\}_{r}^{noise}$.
Here we give an example: When $r=0.00$, the $\{1\}_{r=0.00}^{noise}$ class is identical to the $\{1\}$ class in Fashion-MNIST, comprising 6000 T-shirt images. As $r$ increases, such as $r=0.01$, the noised class $\{1\}_{r=0.01}^{noise}$ still consists of 6000 images, but only 99% of them are T-shirt images, with the remaining 1% randomly selected from other classes. Consequently, when $r=0.9$, $\{1\}_{r=0.9}^{noise}$ contains 600 T-shirt images and 5400 randomly selected images from other classes, i.e., totally randomly selected images in the Fashion-MNIST dataset. The task is finding a single convering polytope that contains $\{1\}_{r=0.9}^{noise}$ against the other data.
We utilize Algorithm 1 to identify a single polytope with minimal width, achieving an accuracy greater than 99.9% on the noised dataset. To mitigate the randomness inherent in image selection, we compute the average value over five repetitions. The results are presented in Table 2.
@@
Table 2: The average number of faces required for a convex polytope to cover the class $\{1\}^{noise}_{r}$, with varying levels of random labels $(r)$. As the proportion of random labels in the class rises, there is a noticeable corresponding increase in the requisite number of faces.
@@
Moreover, our result can demonstrate some previous works that classified these datasets using convex polytopes. Notably, for MNIST (Kantchelian et al., 2014) and CIFAR10 (Sivaprasad et al., 2021), it has been observed that convex polytope classifiers exhibit self-regularization effects and robustness to label noise. Our results presented in Table 2 validate this observation based on the geometric features of the datasets: each class in the datasets inherently possesses simple geometry, enabling convex classifiers to generalize effectively.
A particularly intriguing finding emerges when comparing the results with 50% corrupted labels ($r=0.5$) to those with fully-mixed labels ($r=0.9$). Notably, for MNIST and Fashion-MNIST datasets, Table 2 suggests that the 50% corrupted dataset is even more challenging to segregate than the fully-mixed counterpart. In other words, a convex polytope with fewer faces must encompass the original images from the uncorrupted class $\{1\}$the ’digit 1 images’ in MNIST or the ’T-shirts images’ in FashionMNIST-, necessitating a larger number of faces for the polytope to effectively segregate them. This observation emphasizes the geometric simplicity of MNIST and Fashion-MNIST datasets, contrasting with the behavior observed in CIFAR10, which does not exhibit this trend.
@@ Appendix C Algorithms for finding polytope-basis covers
In this section, we provide detailed explanations of our algorithms, including each step, motivation, and theoretical results. Moreover, we further introduce two algorithms that reduce polytope-basis covers. The section is structured into four subsections.
In Section C.1, we delve into the motivation behind Algorithm 1 and provide a comprehensive explanation of its implementation. In Section C.2, we introduce Algorithm 3, which extracts a polytope-basis cover from a trained two-layer ReLU network. In Section C.3, we unveil Algorithm 4, an efficient algorithm designed to produce a polytope-basis cover for a given dataset using a minimal number of polytopes. Lastly, in Section C.4, we conduct a comparative analysis of our proposed algorithms.
@@ C.1 A polytope-basis cover derived from a three-layer ReLU network
In Section 3.2, we introduced the compressing algorithm (Algorithm 1) that deforms two-layer network to represent a simple convex polytope. Before we provide detail descriptions, we introduce some terminologies.
@@
Roughly speaking, the activation boundaries of $\mathcal{T}$ are non-differentiable points of $\mathcal{T}$. For example, the leftmost column in Figure 2 or 8 shows the decision boundary and activation boundaries of a trained two-layer ReLU network.
Let $\mathcal{T}$ be a two-layer ReLU network defined in (5), thus $v_{k}<0$ for all $k\in[0pt]$. Let $S$ be the region defined by $S:=\{{\bm{x}}~{}|~{}\mathcal{T}({\bm{x}})=1\}$. If it is nonempty, then it is a convex polytope with $0pt$ faces by Definition 2.1). However, its decision boundary $R:=\{{\bm{x}}~{}|~{}\mathcal{T}({\bm{x}})>0\}\supset S$ may contain more data points than $S$. For the given dataset ${{\mathcal{D}}}$, to achieve (7), the objective of the compressing algorithm is to achieve
@@
Note that (9) directly implies (7). Below, we demonstrate a detailed examination of the compressing algorithm, which was briefly introduced in Section 3. The algorithm comprises two parts: 1. eliminating a redundant neuron, and 2. scaling some neurons.
  • PART 1. Eliminating a redundant neuron. The first part involves the removal of remaining redundant neurons that may activate some data points but do not contribute significantly to the network output. Specifically, in Figure 3(a), the neuron indicated by the red arrow does not contribution to the change of the decision boundary since the training data points activated by this neuron already exhibit negative outputs, i.e., $\mathcal{T}({\bm{x}})<0$ already. In essence, eliminating such neurons does not significantly alter the decision boundary of $\mathcal{T}$. However, although the removal of this neuron maintains the decision boundary of $\mathcal{T}$, it does affect the output value of $\mathcal{T}$, consequently influencing other subnetworks in the three-layer network $\mathcal{N}$. To address this, the algorithm removes only one neuron at once, which has the smallest value of $|v_{k}|\cdot\left\|{\bm{w}}_{k}\right\|$.
  • PART 2. Scaling neurons. The second part is deforming the given network to satisfy (9) by magnifying neurons. From the definition of $\mathcal{T}({\bm{x}})=1+\sum_{k}v_{k}\sigma({\bm{w}}_{k}^{\top}{\bm{x}}+b_{k})$, recall that all $v_{k}<0$. If we take $v_{k}\rightarrow-\infty$ for all $k$, then the region $R=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}0<\mathcal{T}({\bm{x}})\}$ shrinks to the region $S=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}({\bm{x}})=1\}$ (cf. Figure 14(b) may help understanding this sentence.) This is the trick we used to figure out the minimal polytope representation of the decision region.
    In this step, we just increase the magnitude of a neuron $(v_{k},{\bm{w}}_{k},b_{k})$ for $k\in[m]$ if it has an activated data ${\bm{x}}_{i}$ such that $\mathcal{T}_{j}({\bm{x}})\neq 1$. The multiplication constant is referred to $\lambda_{scale}$ to $(v_{k},{\bm{w}}_{k},b_{k})$ in the algorithm. Furthermore, note also that the updated neuron still satisfies (8), thus gradient desecent algorithm can be parallelized with keeping $v_{k}<0$. In the example in Figure 3, check that the decision boundary in Figure 3(b) is indeed shrunk in (c).
@@
Figure 2 illustrates the results obtained by applying Algorithm 2. A three-layer network $\mathcal{N}$ defined in (6) with $J=16$ polytopes, each consisting of 20 neurons (architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}320\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}16\rightarrow 1$), was pre-trained on the swiss-roll dataset. Subsequently, Algorithm 1 was applied to compress each $\mathcal{T}_{j}$ with fine tuning, resulting in a compressed three-layer network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}72\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}16\rightarrow 1$. Then, for the polytopes defined by $C_{j}:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}_{j}({\bm{x}})=1\}$, the collection of polytoeps $\mathcal{C}=\{C_{j}\}_{j\in[J]}$ becomes a polytope-basis cover of the dataset, comprising these 16 polytopes of $\mathcal{N}$. The obtained polytoeps are illustrated in Figure 2(b, c).
Furthermore, it is noteworthy to mention the time efficiency of both Algorithm 1 and 2. Despite the presence of multiple for loops, these algorithms are implemented efficiently using parallel computing in PyTorch. Empirically, they demonstrate quick performance with practical neural network widths. On a theoretical level, Proposition C.2 ensures that Algorithm 2 terminates within a finite timeframe and provides a polytope-basis cover that achieved the same accuracy with the converged neural network $\mathcal{N}$.
C.2 A polytope-basis cover derived from a two-layer ReLU network
Figure 8: A polytope-basis cover derived by Algorithm 3 from a trained two-layer ReLU network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}50\rightarrow 1$. The decision boundary and all activation boundaries of the converged network are described in the leftmost column. The algorithm provides a polytope-basis cover consists of 68 polytopes, and some of them are illustrated in other columns.
In this section, we propose another algorithm that extracts a polytope-basis cover from a trained two-layer ReLU network $\mathcal{N}$ defined in (1). First, we decompose $\mathcal{N}$ as the difference of two convex functions:

### v2

@@ 5 Conclusion
In this paper, we investigated the geometric characteristics of datasets and neural network architecture. Specifically, we established both upper and lower bounds on the necessary widths of network architectures for classifying given data manifolds, relying on its polytope-basis cover. Furthermore, we extended these insights to simplicial complexes or spaces consisting of prismatic polytopes, shedding light on how the width bound varies in response to the complexity of the dataset. Conversely, we also demonstrated that the polytope structure of datasets can be inspected by training neural networks. We proposed such an algorithm, and our experimental results unveiled that each class within MNIST, Fashion-MNIST, and CIFAR10 datasets can be distinguished by at most two polytopes, implying geometric simplicity of real-world datasets. We further analyzed that the number of faces in the polytope serves as an indicator of the geometric complexity of each class. Our empirical investigations unveil that neural network indeed converges to a polytope-basis cover of dataset, and conversely, it is possible to inspect geometrical features of the dataset from trained neural networks.
@@
The optimality of the polytope-basis cover obtained by Algorithm 1 has not been clarified, which remains an interesting avenue for future research. Furthermore, while we only considered fully-connected networks in this work, it is desired to investigate the geometry in other network architectures like convolutional neural networks (CNNs).
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT, Ministry of Science and ICT) (No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation), by National Research Foundation of Korea(NRF) (**RS-2023-00262527**), by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST)), and by the National Research Foundation of Korea under Grant RS-2024-00336454
Impact Statement
This paper presents work on describing the polytope structure in deep ReLU networks. This framework provides insights into the geometric roles of neurons and layers, potentially leading to a deeper understanding of high-dimensional dataset geometry and polytope structure of deep ReLU networks. The proposed approach may contribute to investigating data representation, the geometry of feature spaces, and understanding LTH. Lastly, there are no potential societal consequences of this work.
@@ Appendix A. Related Works
Appendix B. Geometric Simplicity of Real-World Datasets
@@ C.2 A Polytope-Basis Cover Derived from a Trained Two-layer ReLU Network
              C.3 An Efficient Algorithm to Find a Simple Polytope-Basis Cover
              C.4 Comparison of the Proposed Algorithms.
@@ Appendix A Related Works
On the other hand, there has been a burgeoning interest in investigating polytope structures induced by deep ReLU networks (Fawzi et al., 2018; Xu et al., 2021; Alfarra et al., 2022; Vincent & Schwager, 2022; Black et al., 2022; Haase et al., 2022; Liu et al., 2023; Fan et al., 2023; Huchette et al., 2023; Vallin et al., 2023; Piwek et al., 2023). Masden (2022) introduced algorithms capable of extracting the polytope structure inherent in networks and deriving topological properties of the decision boundary. Carlsson (2019) and Vallin et al. (2023) considered the pre-image of ReLU networks, characterizing the geometric shapes of the decision boundary to gain an understanding of the polytope partitions of deep ReLU networks.
@@
Recent years have witnessed a surge in research dedicated to unraveling the convexity of ReLU networks concerning their input. Crucially, it has been established that the ReLU activation function demonstrates convexity concerning its input when composited weights are all positive (Proposition 1 in (Amos et al., 2017), cf. Lemma F.4). This discovery led to the inception of Input Convex Neural Networks (ICNNs) by Amos et al. (2017), a characteristic that has been effectively leveraged in various applications (Makkuva et al., 2020; Chen et al., 2020; Fan et al., 2021; Bünning et al., 2021; Balestriero et al., 2022). Notably, Balestriero et al. (2022) applied this idea and proposed DeepHull, the algorithm to approximate the convex hull by convex deep networks. On the other hand, Sivaprasad et al. (2021) asserted that these convex neural networks exhibit self-regularization effects, demonstrating superior generalization performance in specific tasks. Additionally, research has delved into representing neural networks as the difference of convex (DC) functions (Sankaranarayanan & Rengaswamy, 2022; Piwek et al., 2023; Beltran-Royo et al., 2024). Sankaranarayanan & Rengaswamy (2022) employed Linear Programming to optimize polyhedral DC functions, capitalizing on the piecewise linearity induced by the ReLU activation function. (Beltran-Royo et al., 2024) showed that DC NNs have an implicit bias avoiding overfitting in 1-D nonlinear regularization. (Sankaranarayanan & Rengaswamy, 2022) proposed a new network architecture called Convex Difference Neural Network (CDiNN), and suggested using convex concave procedure for optimization. These findings of previous studies suggest that investigating neural networks through the lens of differences in convex functions holds significant potential for future research.
@@
Several studies have delved into the intricate relationship between the geometric features of datasets and neural network training, often referred to as the multiple manifold problem (Goldt et al., 2020; Buchanan et al., 2020; Wang et al., 2021; Chen et al., 2022; Tiwari & Konidaris, 2022). This problem typically involves the binary classification of two low-dimensional manifolds by neural networks. For instance, Buchanan et al. (2020) and Wang et al. (2021) focused on the task of distinguishing between two curves (i.e., one-dimensional manifolds), investigating the convergence speed and generalization concerning the geometric features of the dataset. Similarly, in the context of low-dimensional data manifolds, Tiwari & Konidaris (2022) examined the effects of data geometry on the complexity of trained neural networks, measuring the distance to the manifold. In a related vein, Dirksen et al. (2022) considered the separation problem with randomly initialized ReLU networks, explicitly linking required widths to weight initialization.
Furthermore, there are other numerous empirical studies that have supported the implicit relationship between network architecture and the geometric complexity or topological structure of the data manifold (Fawzi et al., 2018; Kim et al., 2020; Cohen et al., 2020; Birdal et al., 2021; Barannikov et al., 2022, 2021; Naitzat et al., 2020; Hajij & Istvan, 2020, 2021; Magai & Ayzenberg, 2022). Additionally, Li et al. (2018) empirically investigated the loss landscape to measure the intrinsic dimension of datasets, which is deeply related to the minimal neural network architecture. These theoretical and empirical findings suggest a high correlation between neural network architecture and the training dataset: a more complicated dataset requires a more complex architecture.
However, a notable gap still exists in the literature when it comes to explicitly constructing a neural network in practice. For example, as introduced in Section 1, it remains unknown which architecture of neural networks can or cannot completely classify a given dataset with explicit construction of neural networks.
@@ Appendix B Geometric Simplicity of Real-World Datasets
In this section, we present empirical results illustrating how the number of faces of covering polytopes reflects the geometric characteristics of a class. Specifically, we examine a classification task under varying levels of noisy labels. We focus on the class $\{1\}$ in each dataset and manipulate the noise levels (denoted as $r$) across values of 0, 0.01, 0.1, 0.25, 0.5, and 0.90 while maintaining the total number of data points in the class. The noisy class with noise level $r$ is denoted by $\{1\}_{r}^{noise}$.
Here we give an example: When $r=0.00$, the $\{1\}_{r=0.00}^{noise}$ class is identical to the $\{1\}$ class in Fashion-MNIST, comprising 6000 T-shirt images. As $r$ increases, such as $r=0.01$, the noised class $\{1\}_{r=0.01}^{noise}$ still consists of 6000 images, but only 99% of them are T-shirt images, with the remaining 1% randomly selected from other classes. Consequently, when $r=0.9$, $\{1\}_{r=0.9}^{noise}$ contains 600 T-shirt images and 5400 randomly selected images from other classes, i.e., totally randomly selected images in the Fashion-MNIST dataset. The task is finding a single covering polytope that contains $\{1\}_{r=0.9}^{noise}$ against the other data.
We utilize Algorithm 1 to identify a single polytope with minimal width, achieving an accuracy greater than 99.9% on the noised dataset  [^5] Note that Tables 1 and 2 employ different accuracy criteria. With a stricter criterion of 99.99% in Table 1, the identified polytopes are likely more precise but potentially require the larger number of faces. Conversely, the looser criterion of 99.9% used in Table 2 may lead to slightly loose polytope, but easier to identify.. To mitigate the randomness inherent in image selection, we compute the average value over five repetitions. The results are presented in Table 2.
@@
Table 2: The average number of faces required for a convex polytope to cover the noisy class $\{1\}^{noise}_{r}$, with varying levels of random labels $(r)$. As the proportion of random labels in the class rises, there is a noticeable corresponding increase in the requisite number of faces.
@@
Moreover, our result can demonstrate some previous works that classified these datasets using convex polytopes. Notably, for MNIST and CIFAR10 datasets, Kantchelian et al. (2014) and Sivaprasad et al. (2021) observed that convex polytope classifiers exhibit self-regularization effects and robustness to label noise. Our results presented in Table 2 validate this observation based on the geometric features of the datasets: each class in the datasets inherently possesses simple geometry, enabling convex classifiers to generalize effectively.
A particularly intriguing finding emerges when comparing the results with 50% corrupted labels ($r=0.5$) to those with fully-mixed labels ($r=0.9$). Notably, for MNIST and Fashion-MNIST datasets, Table 2 suggests that the 50% corrupted dataset is even more challenging to segregate than the fully-mixed counterpart. In other words, a convex polytope with fewer faces must encompass the original images from the uncorrupted class $\{1\}$the ’digit 1 images’ in MNIST or the ’Trouser images’ in Fashion-MNIST-, necessitating a larger number of faces for the polytope to effectively segregate them. This observation emphasizes the geometric simplicity of MNIST and Fashion-MNIST datasets, contrasting with the behavior observed in CIFAR10, which does not exhibit this trend.
@@ Appendix C Algorithms for Finding Polytope-basis Covers
This section delves into the inner workings of our algorithms. Each algorithm is presented with a breakdown of its steps, the rationale behind its design, and its theoretical underpinnings. Additionally, two novel algorithms dedicated to polytope-basis cover reduction are introduced. For clarity, the section is organized into four sections.
In Section C.1, we delve into the motivation behind Algorithm 1 and provide a comprehensive explanation of its implementation. We introduce Algorithm 3 in Section C.2, which extracts a polytope-basis cover from any trained two-layer ReLU network. Section C.3 introduces Algorithm 4, an efficient algorithm for generating a minimal polytope-basis cover for a given dataset. Finally, Section C.4 presents a comparative analysis of all proposed algorithms.
@@ C.1 A Polytope-Basis Cover Derived From a Three-Layer ReLU Network
In Section 3.2, we introduced the compressing algorithm (Algorithm 1) that deforms a two-layer network to represent a single convex polytope. Before we provide detail descriptions, we introduce some terminologies.
@@
Roughly speaking, the activation boundaries of $\mathcal{T}$ are non-differentiable points of $\mathcal{T}$. For example, the leftmost column in Figure 3 or 9 shows the decision boundary and activation boundaries of trained networks.
Now, we provide in-depth explanations of the algorithms introduced in the main text. Let $\mathcal{T}$ be a two-layer ReLU network defined in (5), thus $v_{k}<0$ for all $k\in[0pt]$. Let $S$ be the region defined by $S:=\{{\bm{x}}~{}|~{}\mathcal{T}({\bm{x}})=\lambda\}$. If it is nonempty, then it is a convex polytope with $0pt$ faces by Definition 2.1. However, its decision boundary $R:=\{{\bm{x}}~{}|~{}\mathcal{T}({\bm{x}})>0\}\supset S$ may contain more data points than $S$. For the given dataset ${{\mathcal{D}}}$, to achieve (7), the objective of the compressing algorithm is to achieve
@@
Note that (9) directly implies (7). Below, we demonstrate a detailed examination of the compressing algorithm step-by-step, which was briefly introduced in Section 3. The algorithm comprises two parts: 1. eliminating a redundant neuron, and 2. scaling some neurons.
  • PART 1. Eliminating a redundant neuron. The first part involves the removal of remaining redundant neurons that may activate some data points but do not contribute significantly to the network output. Specifically, in Figure 3(a), the neuron indicated by the red arrow does not contribute to the change of the decision boundary since the training data points activated by this neuron already exhibit negative outputs, i.e., $\mathcal{T}({\bm{x}})<0$ already. In essence, eliminating such neurons does not significantly alter the decision boundary of $\mathcal{T}$. However, although the removal of this neuron maintains the decision boundary of $\mathcal{T}$, it does affect the output value of $\mathcal{T}$, consequently influencing other subnetworks in the three-layer network $\mathcal{N}$. To address this, the algorithm removes only one neuron at once, which has the smallest value of $|v_{k}|\cdot\left\|{\bm{w}}_{k}\right\|$.
  • PART 2. Scaling neurons. The second part is deforming the given network to satisfy (9) by magnifying neurons. From the definition of $\mathcal{T}({\bm{x}})=\lambda+\sum_{k}v_{k}\sigma({\bm{w}}_{k}^{\top}{\bm{x}}+b_{k})$, recall that all $v_{k}<0$. If we take $v_{k}\rightarrow-\infty$ for all $k$, then the region $R=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}0<\mathcal{T}({\bm{x}})\}$ shrinks to the region $S=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}({\bm{x}})=\lambda\}$ (cf. Figure 15(b)). This is the trick we used to figure out the minimal polytope representation of the decision region.
    In this step, we just increase the magnitude of a neuron $(v_{k},{\bm{w}}_{k},b_{k})$ for $k\in[m]$ if it has an activated data ${\bm{x}}_{i}$ such that $\mathcal{T}_{j}({\bm{x}})\neq\lambda$. The multiplication constant is referred to $\lambda_{scale}$ to $(v_{k},{\bm{w}}_{k},b_{k})$ in the algorithm. Furthermore, note also that the updated neuron still satisfies (8), thus gradient desecent algorithm can be parallelized with keeping $v_{k}<0$. It is noteworthy that in Figure 3(b), the decision boundary demonstrably shrinks to the polytope shown in (c).
@@
Figure 3 illustrates the results obtained by applying Algorithm 2. A three-layer network $\mathcal{N}$ defined in (6) with $J=16$ polytopes, each consisting of 20 neurons (architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}320\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}16\rightarrow 1$), was pre-trained on the swiss-roll dataset. Subsequently, Algorithm 1 was applied to compress each $\mathcal{T}_{j}$ with fine tuning, resulting in a compressed three-layer network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}72\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}16\rightarrow 1$. If the obtained network still completely classify the dataset, then we can derive the polytope basis cover by Theorem 3.4. More precisely, the polytopes defined by $C_{j}:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}_{j}({\bm{x}})=\lambda\}$, the collection of polytoeps $\mathcal{C}=\{C_{j}\}_{j\in[J]}$ becomes a polytope-basis cover of the dataset, comprising these 16 polytopes of $\mathcal{N}$. The obtained polytopes are illustrated in Figure 3(b, c).
Furthermore, it is noteworthy to mention the time efficiency of both Algorithm 1 and 2. Despite the presence of multiple for loops, these algorithms are implemented efficiently using parallel computing in PyTorch. Empirically, they demonstrate quick performance with practical neural network widths. On a theoretical level, Proposition C.2 ensures that Algorithm 2 terminates within a finite timeframe and provides a polytope-basis cover that maintaining the accuracy of the converged neural network $\mathcal{N}$.
C.2 A Polytope-Basis Covers Derived From a Two-Layer ReLU Network
Figure 9: A polytope-basis cover derived by Algorithm 3 from a trained two-layer ReLU network with architecture $2\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}50\rightarrow 1$. The decision boundary and all activation boundaries of the converged network are depicted in the leftmost column. The algorithm provides a polytope-basis cover consists of 68 polytopes, and some of them are illustrated in other columns.
In this section, we propose another algorithm that extracts a polytope-basis cover from a trained two-layer ReLU network $\mathcal{N}$ defined in (1). First, we decompose $\mathcal{N}$ as the sum of convex and concave functions by aligning it according to the sign of the weight values.